Following are the details of sales ( in Lakhs) done by A,B and C. Test for the significant difference in their performance.

             Sales Men

Monthly Sales

    A

48

49

50

49

    B

47

49

48

48

    C

49

51

50

50
Answer #1

one way Anova-

treatment A B C
count, ni = 4 4 4
mean , x̅ i = 49.000 48.00 50.000
std. dev., si = 0.816 0.816 0.816
sample variances, si^2 = 0.667 0.667 0.667
total sum 196 192 200 588 (grand sum)
grand mean , x̅̅ = Σni*x̅i/Σni =   49.00
square of deviation of sample mean from grand mean,( x̅ - x̅̅)² 0.000 1.000 1.000
TOTAL
SS(between)= SSB = Σn( x̅ - x̅̅)² = 0.000 4.000 4.000 8
SS(within ) = SSW = Σ(n-1)s² = 2.000 2.000 2.000 6.000

no. of treatment , k =   3
df between = k-1 =    2
N = Σn =   12
df within = N-k =   9
  
mean square between groups , MSB = SSB/k-1 =    4.0000
  
mean square within groups , MSW = SSW/N-k =    0.6667
  
F-stat = MSB/MSW =    6.0000

anova table
SS df MS F p-value F-critical
Between: 8.00 2 4.00 6.00 0.0221 4.256
Within: 6.00 9 0.667
Total: 14.00 11
α = 0.05

Ho: µ1=µ2=µ3
H1: not all means are equal

F stat = 6.00

Decision:   p-value<α , reject null hypothesis    

conclusion :    there is enough evidence of significance difference among three treatments